OpenStudy (anonymous):

The maximum speed of the Tupolev Tu -144 airliner is 694 m/s. What is this speed in kilometers per hour?

OpenStudy (anonymous):

If you multiply the top and bottom of 694 m/s by 3600, to convert seconds to hours, you get that: 694 m/s = 2498400 m / 3600 s or 694 m/s = 2498400 m/hr You can divide meters by 1000 to get kilometers. 2498400 m/hr = 2498.4 km/hr

OpenStudy (anonymous):

Ah okay, Thank you :)